x represent the side length of a square .find a regular polygon with side length x whose perimeter is twice perimeter of the square. find a regular polygon with side length x whose perimeter is three times the length of the square
Answer by ankor@dixie-net.com(22740)

x represent side length of a square .find a regular polygon with side length x
whose perimeter is twice the perimeter of the square.
:
4x = perimeter of original square
then
8x = twice the perimeter
:
let y = side of polygon with twice the perimeter
2x + 2y = 8x
Simplify, divide by 2
x + y = 4x
y = 4x - x
y = 3x
therefore
2x + 2(3x) = 8x
Polygon with twice the perimeter: 3x by x
:
find a regular polygon with side length x whose perimeter is three times
the length of the square
this perimeter: 3(4x) = 12x
let y = side of polygon with three times the perimeter
2x + 2y = 12x
Simplify, divide by 2
x + y = 6x
y = 6x - x
y = 5x
therefore
2x + 2(5x) = 12x
Polygon with three times the perimeter: 5x by x